# Token Reduction Should Go Beyond Efficiency in Generative Models – From Vision, Language to Multimodality

### v1

@@ Abstract
In Transformer architectures, tokens—discrete units derived from raw data—are formed by segmenting inputs into fixed-length chunks. Each token is then mapped to an embedding, enabling parallel attention computations while preserving the input’s essential information. Due to the quadratic computational complexity of transformer self-attention mechanisms, token reduction has primarily been used as an efficiency strategy. This is especially true in single vision and language domains, where it helps balance computational costs, memory usage, and inference latency. Despite these advances, this paper argues that token reduction should transcend its traditional efficiency-oriented role in the era of large generative models. Instead, we position it as a fundamental principle in generative modeling, critically influencing both model architecture and broader applications. Specifically, we contend that across vision, language, and multimodal systems, token reduction can: (i) facilitate deeper multimodal integration and alignment, (ii) mitigate "overthinking" and hallucinations, (iii) maintain coherence over long inputs, and (iv) enhance training stability, etc. We reframe token reduction as more than an efficiency measure. By doing so, we outline promising future directions, including algorithm design, reinforcement learning-guided token reduction, token optimization for in-context learning, and broader ML and scientific domains. We highlight its potential to drive new model architectures and learning strategies that improve robustness, increase interpretability, and better align with the objectives of generative modeling.  [^1] We collected a list of token reduction papers at Github.
@@ 2.1 Token Reduction in Vision Models
Image Classification. Classification serves as a fundamental task for vision models and token reduction techniques have been widely applied in it due to its simplicity and versatility. It has been widely explored from various aspects bolya2022token ; liang2022not ; rao2021dynamicvit ; wu2023ppt ; zeng2024m2m . Specifically, DynamicViT rao2021dynamicvit devises a lightweight module to predict the importance score of each token, thereby pruning unimportant tokens. E-ViT liang2022not identifies attentive tokens from feed-forward network, enabling token pruning without additional parameters. ToMe bolya2022token merges tokens with similarity based on bipartite matching to maintain information utility. PPT wu2023ppt analyzes the statistic data between layers and adaptively employs token pruning and merging within layers to achieve higher acceleration performance.
Year2020 - 2021Power-bert goyal2020power , DynamicViT rao2021dynamicvit ,TokenLearner ryoo2021tokenlearner , SpAtten wang2021spatten 2022LTP kim2022learned , SPViT kong2022spvit , EViT liang2022not ,PatchSlim tang2022patch , STTS wang2022efficient , PPT wu2023ppt 2023DCP anagnostidis2023dynamic , ToMe bolya2022token , ToMeSD bolya2023token , PuMer cao2023pumer ,HeatViT dong2023heatvit , Llmlingua jiang2023llmlingua ,DTP nawrot2022efficient , TPS wei2023joint 2024Honeybee cha2024honeybee , FastV chen2024image , TokenPacker li2024tokenpacker ,BRAVE kar2024brave , LLaMA-VID li2024llama , DyDiT zhao2024dynamic - 2025.05DivPrune alvar2025divprune , SeTok wu2024towards , TopV yang2025topv ,Voco-llama ye2024voco , S2-MAD zeng2025s , ToCa zou2025accelerating Vision & Language TransformerLLM & Multimodal LLMEfficiency
@@
$\bm{\Rightarrow}$  Future
Figure 1: Timeline of notable method developments for token reduction methods with modality shifts (Vision & Language $\to$ Multimodal LLMs). All these strategies aim to speed up inference with neglected performance drops. Conversely, we ask: What is the next token reduction paradigm in generative model design that goes beyond test-time accelerations?
Video Compression. Unlike token reduction in image classification, video compression focuses more on the temporal redundancy within videos, and algorithms are developed to reduce the number of tokens with less computational overhead. Various token reduction methods have been investigated for different tasks including video understanding ryoo2021tokenlearner ; wang2022efficient , video editing li2024vidtome , video-text retrieval liu2022ts2 ; shen2024tempme , video action detection chen2023efficient and so on. Specifically, STTS wang2022efficient introduces a lightweight framework that dynamically selects the most informative spatial-temporal tokens in video transformers. Tokenlearner ryoo2021tokenlearner proposes an adaptive tokenization module that learns a handful of informative spatial-temporal tokens, significantly reducing computational costs. EVAD chen2023efficient selectively drops irrelevant spatial-temporal tokens in non-keyframes while preserving keyframe and motion-relevant tokens, and then refines actor features using a context-aware decoder to maintain accuracy with reduced computations.
@@ 5.2 Token Reduction for Reinforcement Learning
Reinforcement learning (RL)-driven token reduction technology demonstrates significant potential in multimodal large models, with its core lying in balancing computational efficiency and reasoning accuracy through dynamic reward mechanisms and sparsity constraints. Under the “Fast-Slow Thinking” framework xiao2025fast , RL can guide hierarchical screening of high-value tokens: The fast path employs sparsity strategies (e.g., rule-based rewards or information density scoring) to filter redundant visual or semantic features, while the slow path focuses on refined reasoning. This hierarchical mechanism not only reduces computational overhead by up to 30-50%, but also continuously optimizes token selection strategies via RL’s online feedback. For instance, integrating self-play mechanisms with the GRPO algorithm (Group Relative Policy Optimization) enables the generation of diverse candidate paths and selects optimal outputs.
@@
Furthermore, the sparsity reinforcement paradigm proposed in ZipR1 chen2025zipr1reinforcingtokensparsity highlights RL’s unique advantage in cross-modal generalization. By designing sparsity rewards (e.g., token quantity limits or step compression incentives), models can autonomously generate concise intermediate reasoning chains and even distill such capabilities into lightweight models. For example, in visual question answering tasks, models can dynamically extract key image region features while enhancing answer reliability through multi-candidate contrast mechanisms (e.g., parallel samplings), achieving the goal of “low computational cost, high task accuracy.” Looking ahead, this technology could advance toward adaptive sparsity and cross-modal alignment, providing efficient inference foundations for edge computing and real-time interaction scenarios. Simultaneously, it will drive MLLMs toward co-evolution of lightweight design and strong robustness.
5.3 From Prompt Tuning to In-context-learning
@@
Current token reduction efforts for prompts have primarily aimed at compressing prompts for efficiency, often with impressive results jiang2023llmlingua ; mu2023learning . Looking forward, token reduction should evolve into enhancing reasoning and maximizing utility per token in context. Instead of focusing solely on making prompts shorter, future research should explore how each remaining token can carry more information or trigger more complex inference during in-context learning. One direction is to alter the generation paradigm itself, for example, training language models to predict multiple tokens per step gloeckle2024better . Another idea is to enable deeper internal reasoning without increasing prompt length. CoD adams2023sparse introduces an iterative approach using token reduction within fixed-length constraints: Each step incorporates new salient entities into the summary by compressing/abstracting existing content, forcing the model to enhance density without expanding length.
@@
In summary, the next phase of token reduction research should shift focus from simple prompt compression to reasoning-centric compression. Rather than just trimming prompts, we should ask: How can we make each token in the prompt or context do more work for us? This involves training models with objectives that reward higher-level inference per token, developing architectures that recycle tokens for multi-step thinking, or dynamically selecting the most salient tokens to keep at each stage of dialogue.
@@ 5.6 Towards Long Video Applications
Exploiting long videos holds great potential, as processing hours of footage is significantly more labor-intensive and time-consuming than working with short clips. Due to the inherent complexity and resource demands, most current research on long video learning focuses on discriminative tasks such as video understanding lee2024video ; ren2023testa . In contrast, broader applications including long video editing zhang2025adaflow , long video-text retrieval, and narrative-level generation remain largely underexplored. Progress in these areas could have a significant impact on scene editing in video clips, character rendering in movies, and retrieving useful information from numerous videos. Moreover, token reduction offers a path toward interpretability and efficiency in long video processing jiang2025token . This mimics the human visual system, which does not attend to every frame in detail but instead focuses on salient spatiotemporal changes, such as actions or object movement, while filtering out static, redundant content like backgrounds and stationary objects. Future models should similarly prioritize informative frames and temporal segments, allowing them to reason over extended video sequences with greater efficiency and interpretability.
@@ 5.7 Towards AI for Broader ML and Scientific Domains
To enhance informativeness and reduce redundancy, MedTok employs a token packing mechanism. It optimizes shared tokens and modality-specific tokens, ensuring that the final tokens encode both shared semantic meaning and modality-specific structure. This process drastically reduces effective vocabulary size, addressing the scalability challenge of 600,000+ medical codes by collapsing redundant representations while preserving critical clinical context. Inspired by adaptive tokenization methods for vision duggal2024adaptive ; yan2024elastictok ; kang2023tictok , future EHR tokenization would be adaptive, enabling dynamic representation of patients’ medical histories, where the length of the token series of each patient history would be directly correlated to length and complexity. Such adaptive tokenization can significantly improve training and inference efficiency across diverse healthcare systems.

### v2

@@ Abstract
In Transformer architectures, tokens—discrete units derived from raw data—are formed by segmenting inputs into fixed-length chunks. Each token is then mapped to an embedding, enabling parallel attention computations while preserving the input’s essential information. Due to the quadratic computational complexity of transformer self-attention mechanisms, token reduction has primarily been used as an efficiency strategy. This is especially true in single vision and language domains, where it helps balance computational costs, memory usage, and inference latency. Despite these advances, this paper argues that token reduction should transcend its traditional efficiency-oriented role in the era of large generative models. Instead, we position it as a fundamental principle in generative modeling, critically influencing both model architecture and broader applications. Specifically, we contend that across vision, language, and multimodal systems, token reduction can: (i) facilitate deeper multimodal integration and alignment, (ii) mitigate "overthinking" and hallucinations, (iii) maintain coherence over long inputs, and (iv) enhance training stability, etc. We reframe token reduction as more than an efficiency measure. By doing so, we outline promising future directions, including algorithm design, reinforcement learning-guided token reduction, token optimization for in-context learning, and broader ML and scientific domains. We highlight its potential to drive new model architectures and learning strategies that improve robustness, increase interpretability, and better align with the objectives of generative modeling.  [^1] We collected a list of token reduction papers at: Awesome-Collection-Token-Reduction.
@@ 2.1 Token Reduction in Vision Models
Image Classification. Classification serves as a fundamental task for vision models and token reduction techniques have been widely applied in it due to its simplicity and versatility. It has been widely explored from various aspects bolya2022token ; liang2022not ; rao2021dynamicvit ; wu2023ppt ; zeng2024m2m ; kong2023peeling ; zeng2025token . Specifically, DynamicViT rao2021dynamicvit devises a lightweight module to predict the importance score of each token, thereby pruning unimportant tokens. SPViT kong2022spvit introduces a soft pruning technique, which integrates the less informative tokens generated by the selector module into a package token that will participate in subsequent calculations rather than being completely discarded. EViT liang2022not identifies attentive tokens from the attention map, enabling token pruning without additional parameters. ToMe bolya2022token merges tokens with similarity based on bipartite matching to maintain information utility. PPT wu2023ppt analyzes the statistic data between layers and adaptively employs token pruning and merging within layers to achieve higher acceleration performance.
Year2020 - 2021Power-bert goyal2020power , DynamicViT rao2021dynamicvit ,TokenLearner ryoo2021tokenlearner , SpAtten wang2021spatten 2022LTP kim2022learned , SPViT kong2022spvit , EViT liang2022not ,PatchSlim tang2022patch , STTS wang2022efficient , PPT ma2022ppt 2023DCP anagnostidis2023dynamic , ToMe bolya2022token , ToMeSD bolya2023token , PuMer cao2023pumer ,HeatViT dong2023heatvit , Llmlingua jiang2023llmlingua ,DTP nawrot2022efficient , TPS wei2023joint 2024Honeybee cha2024honeybee , FastV chen2024image , TokenPacker li2024tokenpacker ,BRAVE kar2024brave , LLaMA-VID li2024llama , DyDiT zhao2024dynamic - 2025.07DivPrune alvar2025divprune , SeTok wu2024towards , TopV yang2025topv , SparseVLM zhang2024sparsevlm Voco-llama ye2024voco , S2-MAD zeng2025s , ToCa zou2025accelerating ,TALE han2024token Vision & Language TransformerLLM & Multimodal LLMEfficiency
@@
$\bm{\Rightarrow}$  Future ?
Figure 1: Timeline of notable method developments for token reduction methods with modality shifts (Vision & Language $\to$ Multimodal LLMs). All these strategies aim to speed up inference with negligible performance drops. Conversely, we ask: What is the next token reduction paradigm in generative model design that goes beyond test-time accelerations?
Video Compression. Unlike token reduction in image classification, video compression focuses more on the temporal redundancy within videos, and algorithms are developed to reduce the number of tokens with less computational overhead. Various token reduction methods have been investigated for different tasks, including video understanding ryoo2021tokenlearner ; sun2025llavascissortokencompressionsemantic ; wang2022efficient , video editing li2024vidtome , video-text retrieval liu2022ts2 ; shen2024tempme , video action detection chen2023efficient , and so on. Specifically, STTS wang2022efficient introduces a lightweight framework that dynamically selects the most informative spatial-temporal tokens in video transformers. Tokenlearner ryoo2021tokenlearner proposes an adaptive tokenization module that learns a handful of informative spatial-temporal tokens, significantly reducing computational costs. EVAD chen2023efficient selectively drops irrelevant spatial-temporal tokens in non-keyframes while preserving keyframe and motion-relevant tokens, and then refines actor features using a context-aware decoder to maintain accuracy with reduced computations.
@@ 5.2 Token Reduction for Reinforcement Learning
Reinforcement learning (RL)-driven token reduction technology demonstrates significant potential in multimodal large models, with its core lying in balancing computational efficiency and reasoning accuracy through dynamic reward mechanisms and sparsity constraints wang2025beyond . Under the “Fast-Slow Thinking” framework xiao2025fast , RL can guide hierarchical screening of high-value tokens: The fast path employs sparsity strategies (e.g., rule-based rewards or information density scoring) to filter redundant visual or semantic features, while the slow path focuses on refined reasoning. This hierarchical mechanism not only reduces computational overhead by up to 30-50%, but also continuously optimizes token selection strategies via RL’s online feedback.
@@
The sparsity reinforcement paradigm proposed in ZipR1 chen2025zipr1reinforcingtokensparsity highlights RL’s unique advantage in cross-modal generalization. By designing sparsity rewards (e.g., token quantity limits or step compression incentives), models can autonomously generate concise intermediate reasoning chains and even distill such capabilities into lightweight models. For example, in visual question answering tasks, models can dynamically extract key image region features while enhancing answer reliability through multi-candidate contrast mechanisms (e.g., parallel samplings), achieving the goal of “low computational cost, high task accuracy.”
Recent findings further reveal that optimizing only a small subset of high-entropy "forking tokens" can match or even surpass full-gradient updates in reasoning tasks wang2025beyond , suggesting that token-level entropy patterns can guide efficient policy updates. This entropy-aware strategy reinforces the broader view that sparse, selective token refinement is sufficient for preserving reasoning performance in large models. Looking ahead, the integration of such approaches could enhance cross-modal alignment and inference efficiency in real-time and resource-constrained scenarios, supporting a new generation of lightweight yet capable multimodal language agents.
5.3 From Prompt Tuning to Chain of Thought Reasoning
@@
Current token reduction efforts for prompts have primarily aimed at compressing prompts for efficiency, often with impressive results jiang2023llmlingua ; mu2023learning . Looking forward, token reduction should evolve into enhancing reasoning and maximizing utility per token in context. Instead of focusing solely on making prompts shorter, future research should explore how each remaining token can carry more information or trigger more complex inference during in-context learning and chain of thought reasoning. One direction is to alter the generation paradigm itself, for example, training language models to predict multiple tokens per step gloeckle2024better . Another idea is to enable deeper internal reasoning without increasing prompt length adams2023sparse .
As mentioned in Sec. 4, long CoT chains can become verbose: excessive reasoning steps may introduce errors or obscure logical clarity, particularly in LLM agents where internal reasoning alternates with external tool use gao2025txagent ; wang2024toolgen . Token reduction may serve a critical role in this context by compressing intermediate reasoning into a compact representation. For example, approaches based on next-token prediction mu2023learning ; zhang2025lightthinker can distill intermediate thinking chains into a set of dense, information-rich tokens. These compressed representations can then replace the full intermediate context and serve as inputs for subsequent reasoning steps.
This compressed-thinking strategy has two main benefits: 1) reducing error accumulation and keeping the logic clear by focusing on key information, and 2) allowing more reasoning rounds to fit within a fixed context window, enabling deeper multi-step inference without exceeding length limits.
@@
In summary, the next phase of token reduction research should shift focus from simple prompt compression to reasoning-centric compression. Rather than just trimming prompts, we should ask: How can we make each token in the prompt or context do more work for us? This involves training models with objectives that reward higher-level inference per token, developing architectures that recycle tokens for multi-step thinking, or dynamically selecting the most salient tokens to keep at each step of reasoning.
@@ 5.6 Towards Long Video Applications
Exploiting long videos holds great potential, as processing hours of footage is significantly more labor-intensive and time-consuming than working with short clips. Due to the inherent complexity and resource demands, most current research on long video learning focuses on discriminative tasks such as video understanding lee2024video ; ren2023testa . In contrast, broader applications including long video editing zhang2025adaflow , long video-text retrieval, and narrative-level generation remain largely underexplored. Progress in these areas could have a significant impact on scene editing in video clips, character rendering in movies, and retrieving useful information from numerous videos.
Moreover, token reduction offers a path toward interpretability and efficiency in long video processing jiang2025token . This mimics the human visual system, which does not attend to every frame in detail but instead focuses on salient spatiotemporal changes, such as actions or object movement, while filtering out static, redundant content like backgrounds and stationary objects. Future models should similarly prioritize informative frames and temporal segments, allowing them to reason over extended video sequences with greater efficiency and interpretability.
@@ 5.7 Towards AI for Broader ML and Scientific Domains
To enhance informativeness and reduce redundancy, MedTok employs a token packing mechanism. It optimizes shared tokens and modality-specific tokens, ensuring that the final tokens encode both shared semantic meaning and modality-specific structure. This process drastically reduces effective vocabulary size, addressing the scalability challenge of 600,000+ medical codes by collapsing redundant representations while preserving critical clinical context. Inspired by adaptive tokenization methods for vision duggal2024adaptive ; kang2023tictok ; yan2024elastictok , future EHR tokenization would be adaptive, enabling the dynamic representation of patients’ medical histories, where the length of the token series for each patient’s history would be directly correlated with the length and complexity. Such adaptive tokenization can significantly improve training and inference efficiency across diverse healthcare systems.